How do I solve this problem.6a2-b2 when a=1/3 and b=1/3
Answer by stanbon(75887)

6a2-b2 when a=1/3 and b=1/3
-----
6(1/3)^2 -(1/3)^2
---
= 5(1/3)^2
---
= 5(1/9)
----
= 5/9
----------
Cheers,
Stan H.
===========